Write the equation of the line with slope=0 and containing the point (8,-2)
Answer by Theo(13342)

y = -2
the value of x can be anything, including 8.
the slope intercept form of the equation would be:
y = mx + b
since the slope is 0, the equation becomes:
y = 0x + b
b is the y intercept.
to find b, substitute a coordinate point on the line for the value of x and y to get:
point = (8,-2)
y = 0x + b becomes:
-2 = 0*8 + b
this becomes:
-2 = b
the y intercept is b and the equation becomes:
y = 0x -2
the 0x cancels out of the equation and you are left with:
y = -2